• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S.Beck

    Log entry time 18:46:20 on August 1,2006

    Entry number 176732

    keyword=Shift summary

    Shift summary:

    8:00 - Keep taking production data. Target: LD2 (radiator IN), current: 12uA. Dead time too high: 29% on HRSL.

    8:30 - 8:40 We get no beam due to some checks in Hall C.

    8:40 - Back to data taking, with a lower current - 11uA, to reduce the dead time.

    8:40-8:43 - Took a short run with many beam trips. (I=11uA, DT=27%)

    8:43 - Beam trips - dipole trips in the Hall. MCC tries to fix it.

    8:50 - Beam is back, with 10uA (to further reduce dead time)


    ---- Run summary ----------
    8:15-8:29: run 1971 - production with LD2, current: 12uA, rate=2.8kHz, DT=29%.

    8:38-8:42: run 1972 - A short run with many beam trips. Current: 11uA, rate=2.6 kHz, DT=27%.

    8:53-9:14: run 1973, curent: 10uA, rate=2.2 kHz, DT=25%.

    9:15- 9:34 run 1974, current: 10uA, rate=2.2 kHz, DT=24%.

    10:00 run stop due to Moller measurement.
    10:00-16:00 Moller